OpenStudy (anonymous):

T(n) is the total number of lights in a triangular lighting rig of n rows, as shown. Which polynomial function represents T(n + 1)? T(n)=1/2n^2+1/2n a. T(n+1)=1/2n^2+3/2n+1 b. T(n+1)=1/2n^2+1/2n+1 c. T(n+1)=1/2n^2+n+1 d. T(n+1)=1/2n^2+1

OpenStudy (blacksteel):

You'd solve this by simply replacing n with (n+1) everywhere in the equation and then simplifying. T(n) = 0.5n^2 + 0.5n T(n+1) = 0.5(n+1)^2 + 0.5(n+1) =0.5(n^2 + 2n + 1) + 0.5(n+1) =0.5n^2 + n + 0.5 + 0.5n + 0.5 =0.5n^2 + 1.5n + 1 or (1/2)n^2 + (3/2)n + 1

OpenStudy (anonymous):

oh i get it now thx